ROTATIONAL SPECTROSCOPY AND RING-PUCKERING CONFORMATION OF 3-HYDROXY-TETRAHYDROFURAN

RICHARD J. LAVRICH, RACHELE L. RHEA, JAMES W. McCARGAR, and \underline{MICHAEL J. TUBERGEN}, Department of Chemistry, Kent State University, Kent, OH 44242.

The ring puckering conformation of 3-hydroxytetrahydrofuran, a five-membered heterocyclic ring, has been determined from rotational spectroscopy. Rotational constants for the normal isotopomer are A = 5581.8230 (7), B = 3638.8316 (7), and C = 2924.7410 (7) MHz; rotational spectra were also recorded for four ^{13}C isotopic species in natural abundance. The best fit structure is a C_{4′} endo conformation with an intramolecular hydrogen bond from the hydroxyl to the ring oxygen. Projections of the dipole moment on the inertial axes, μ_{a} = 0.956 (1) D, μ_{b} = 0.875 (2) D, and μ_{c} = 1.050 (2) D, were determined from Stark effect measurements. The experimental structure is nearly identical to the lowest energy ab initio conformation (MP2/6-31+G**).